On May 1, 2018, Karina purchased 100 shares of Gold stock at a total cost of $2,000. She received a total of $150 in dividends and sold the stock for $2,280 on May 15, 2019. Karina has a combined state and federal marginal tax rate of 28%. Her tax rate on both long-term capital gains and dividend income is 15%. What is her after-tax holding period return on her investment in Gold stock?

21.50%

23.48%

18.28%

16.28%
Answer #1

Purchase price of stock = $2,000

Sell price of Stock = $2,280

Dividend received = $150

Tax rate on capital gain and dividend income = 15%

Total profit during holding = (2280-2000) + 150

= $430

Total profit after tax = 430*(1-0.15)

= $365.5

Thus,

After tax Holding period Return = Total profit after tax/purchase price

= 365.5/2000

= 0.18275

= 18.28 %

Hope this will help, please do comment if you need any further explanation. Your feedback would be appreciated.